OpenStudy (anonymous):

Eliminate the parameter. x = 4t - 4 y = 16t2 - 1 A. y = x2 + 8x + 15 B. y = 16x2 - 8x + 3 C. y = x2 + 8x + 17 D. y = 16x2 - 8x + 5

OpenStudy (luigi0210):

Start by setting the first equation equal to t

OpenStudy (anonymous):

t=-x(4)+4 ?

OpenStudy (luigi0210):

x=4t-4 x+4=4t (x/4)+1=t

OpenStudy (anonymous):

and then after that you plug in to t?

OpenStudy (luigi0210):

Yup, plug that in to the second equation

OpenStudy (dan815):